OpenStudy (anonymous):

i need help in this problem:

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{a-x}=\sqrt{y}-\sqrt{a-y}=\sqrt{z}-\sqrt{a-z}\] prove that (y-z)(z-x)(x-y)=0

OpenStudy (jamesj):

Is this top equation right? should it be sqrt(x) MINUS sqrt(a-x)?

OpenStudy (anonymous):

oh yeah, sorry

OpenStudy (anonymous):

Square them up \[a - 2 *\sqrt{x} * \sqrt{a-x} = a - 2*\sqrt{y}*\sqrt{a-y} = a - 2*\sqrt{z}*\sqrt{a-z}\] \[\sqrt{x} * \sqrt{a-x}=\sqrt{y}*\sqrt{a-y}=\sqrt{z}*\sqrt{a-z}\]

OpenStudy (anonymous):

next??

OpenStudy (anonymous):

\[x(a-x)=y(a-y)=z(a-z)\]

OpenStudy (anonymous):

then??

OpenStudy (anonymous):

Now they all are equal right! so I think any two of them should be same I mean either x=y or y=z or x=z

OpenStudy (anonymous):

(y-z)(z-x)(x-y)=0

OpenStudy (anonymous):

squaring each expression x - 2sqrt(a-x) + a - x = -2sqrt(a-x) + a also we get -2sqrt(a-y) + a - 2sqrt(a-z) + a these 3 are equal so (a-y)=(a-x)=(a-z) so x=y=z there fore (y-z)(z-x)(x-y)=0

OpenStudy (jamesj):

@jimmy, you've dropped sqrt(x) etc.

OpenStudy (anonymous):

there should not b sqrt{x(a-x)}??

OpenStudy (anonymous):

@jamesj, thats my point

OpenStudy (anonymous):

\[x(a-x)=y(a-y)=z(a-z)\] \[xa - x^2 = ya - y^2 \] \[ay -y^2 = az -z^2\] \[xa - x^2 = az -z^2 \] x,y,z can have two values but not three so eventually for the equality two variables must be same

OpenStudy (anonymous):

or if you want to solve it more you will see x +y = a z+y=a x+z=a

OpenStudy (anonymous):

yes - i see - right james - should be 2[sqrtx* sqrt(a-x)] etc

OpenStudy (jamesj):

Starting with Ishaan's x(a−x)=y(a−y) we have x^2 - ax + y(a-y) = 0 =>\[x = \frac{1}{2} \left( a \pm \sqrt{a^2 -4ya+4y^2} \right) = \frac{1}{2} (a \pm (a-2y) ) =y \ \ or \ \ a\]

OpenStudy (jamesj):

If x = y we're finished as x - y = 0 If x = a then y = 0 or y = a If y = 0 and x = a, then z = 0 or a and we're finished as (x-y) or (y-z) = 0 If y = a and x = a, then x - y = 0

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

thanx, guys

OpenStudy (jamesj):

Sorry x =y or a - y. Now modify my next argument, but it still works.

OpenStudy (phi):

Or, (to beat a dead horse), starting with x +y = a z+y=a x+z=a subtract equations: (x+y)-(z+y)= 0 or (x-z)=0 and so on to get all 3 terms